OpenStudy (anonymous):

(x - 3)(6x - 2) find product :)

OpenStudy (anonymous):

\[(x - 3)(6x - 2)\] \[x \times 6x + x \times (-2) + (-3) \times (6x) + (-3) \times (-2)\] \[6x^2 -2x -18x +6\] \[6x^2 -20x +6\]